OpenStudy (anonymous):

A poll was conducted in states A, B, and C to determine which candidate voters would most likely vote for. The results are displayed in the table below: State A State B State C Total Liberal 17 20 12 49 Conservative 23 15 13 51 Total 40 35 25 100 What is P(conservative | State B)? 0.35 0.43 0.49 0.57

OpenStudy (kirbykirby):

\[P(\text{conservative }|\text{ state B})=\frac{P(\text{conservative } \cap \text{ state B})}{P(\text{state B})}=\frac{15/100}{35/100}=3/7=0.43 \]